A BOAT IS MOVING WITH VELOCITY 30 M/S TOWARDS EAST. AFTER 10S ITS VELOCITY BECOMES 40M/S TOWARDS NORTH. THE AVERAGE ACCELERATION OF THE BODY IS
 

Grade:11

4 Answers

Shailendra Kumar Sharma
6 years ago
Assume average acceleration concept = (V2-V1)/time=(40j-30i)/10=4j-3iSo magnitude of acceleration is 5
Arun
6 years ago
Tr
Arun
6 years ago
Initial velocity = The velocity in east direction = 30i , where i is unit vector along x axis.
Final velocity = The velocity in north direction = 40j , where j is unit vector along y axis. 
Average acceleration = { (root over(302+402))/10} m/s2 
                               = 5 m/s2 in north east direction! 
Lakshay Kumar
6 years ago
Initial velocity, u=30iFinal velocity, v=40jNow apply v=u+at40j=30i+a(10)10a=40j-30ia=4j-3i=>a=5m/s^2 in North-west direction